Barton Industries estimates its cost of common equity by using three approaches: the CAPM, the bond-yield-plus-risk-premium approach, and the DCF model. Barton expects next year's annual dividend, D1, to be $2.20 and it expects dividends to grow at a constant rate g = 4.4%. The firm's current common stock price, P0, is $20.00. The current risk-free rate, rRF, = 4.7%; the market risk premium, RPM, = 6.2%, and the firm's stock has a current beta, b, = 1.35. Assume that the firm's cost of debt, rd, is 14.53%. The firm uses a 3.2% risk premium when arriving at a ballpark estimate of its cost of equity using the bond-yield-plus-risk-premium approach. What is the firm's cost of equity using each of these three approaches? Round your answers to two decimal places.

CAPM cost of equity: %

Bond yield plus risk premium: %

DCF cost of equity: %

What is your best estimate of the firm's cost of equity? -Select-
Answer #1

1.
=4.7%+1.35*6.2%=13.07%

2.
=14.53%+3.2%=17.73%

3.
=2.2/20+4.4%=15.40%

4.
=1/3*(4.7%+1.35*6.2%+14.53%+3.2%+2.2/20+4.4%)=15.40%